A soccer field is 400 ft long and 150 ft wide. What is the field’s dimension in meters and its area in square centimeters?



1
Expert's answer
2021-08-30T06:54:23-0400

(a)

"Length=400\ ft\cdot\dfrac{0.3048\ m}{1\ ft}=121.92\ m,""Width=150\ ft\cdot\dfrac{0.3048\ m}{1\ ft}=45.72\ m."

(b)

"Area=Length\cdot Width,""Area=400\ ft\cdot150\ ft=6\cdot10^4\ ft^2\cdot(\dfrac{30.48\ cm}{1\ ft})^2,""Area=5.6\cdot10^7\ cm^2."